# Charged Nariai black holes on the dark bubble

### v1

@@ Appendix A The dilatonic coupling
$S_{\text{NSNS}}=\frac{1}{2\kappa_{10}^{2}g_{s}^{2}}\int\mathrm{d}x^{10}\sqrt{-g_{10}}\left(R_{10}-\frac{1}{2}(\partial\phi)^{2}-\frac{1}{2}\mathrm{e}^{-(\phi-\phi_{0})}|H|^{2}\right)\,,$ (99)
@@
$S_{\text{NSNS}}=\frac{1}{2\kappa_{5}^{2}}\int\mathrm{d}x^{5}\sqrt{-g_{5}}\left(R_{5}-\frac{1}{2}(\partial\phi)^{2}-\frac{1}{2}\mathrm{e}^{-(\phi-\phi_{0})}|H|^{2}\right)\,,$ (100)
@@
Notice that a 5d cosmological constant can arise from moduli stabilisation from internal fluxes (RR and NSNS). We also remark that there is a dilatonic coupling to the $H$-field. We have not included it in the main text, as we assumed the dilaton to be constant, such that the dilatonic coupling disappears. We justify this assumption here below. With the presence of a dynamical dilaton, we infer that its equation of motion is
@@
which justifies taking the dilaton to be constant.
@@ Appendix B Energy-momentum tensor on the brane
$S_{\text{DBI}}=-\sigma\int\mathrm{d}x^{4}\sqrt{-\det\left(h_{ab}+\tau\mathcal{F}_{ab}\mathrm{e}^{-(\phi-\phi_{0})/2}\right)}\,,$ (105)
@@
$S_{\text{DBI}}=-\sigma\int\mathrm{d}x^{4}\sqrt{-\det\left(h_{ab}+\tau\mathcal{F}_{ab}\right)}\,.$ (107)
@@
$S_{\text{DBI}}=\int\mathrm{d}x^{4}\sqrt{-h_{4}}\left(-\sigma-\frac{\sigma\tau^{2}}{2}|\mathcal{F}|^{2}\right)$ (108)
This reminds of the Maxwell action where $\sigma\tau^{2}$ is the inverse gauge coupling squared. This field strength sources the $H$-field. Indeed, we find that

### v2

@@ Appendix A The dilatonic coupling
$S_{\text{NSNS}}=\frac{1}{2\kappa_{10}^{2}g_{s}^{2}}\int\mathrm{d}^{10}x\sqrt{-g_{10}}\left(R_{10}-\frac{1}{2}(\partial\phi)^{2}-\frac{1}{2}\mathrm{e}^{-(\phi-\phi_{0})}|H|^{2}\right)\,,$ (99)
@@
$S_{\text{NSNS}}=\frac{1}{2\kappa_{5}^{2}}\int\mathrm{d}^{5}x\sqrt{-g_{5}}\left(R_{5}-\frac{1}{2}(\partial\phi)^{2}-\frac{1}{2}\mathrm{e}^{-(\phi-\phi_{0})}|H|^{2}\right)\,,$ (100)
@@
Notice that a 5d cosmological constant can arise from moduli stabilisation from internal fluxes (RR and NSNS). We also remark that there is a dilatonic coupling to the $H$-field. We have not included it in the main text, as we assumed the dilaton to be constant, so the dilatonic coupling disappears. We justify this assumption here below. With the presence of a dynamical dilaton, we infer that its equation of motion is
@@
which justifies taking the dilaton to be constant. If one wants to solve our setup beyond this order, it is necessary to keep dilatonic couplings in the 5d Einstein equations and solve for all the metric functions and dilaton accordingly, essentially solving for a dilatonic, magnetic AdS5 black string. This is beyond the scope of this work.
@@ Appendix B Energy-momentum tensor on the brane
$S_{\text{DBI}}=-\sigma\int\mathrm{d}^{4}x\sqrt{-\det\left(h_{ab}+\tau\mathcal{F}_{ab}\mathrm{e}^{-(\phi-\phi_{0})/2}\right)}\,,$ (105)
@@
$S_{\text{DBI}}=-\sigma\int\mathrm{d}^{4}x\sqrt{-\det\left(h_{ab}+\tau\mathcal{F}_{ab}\right)}\,.$ (107)
@@
$S_{\text{DBI}}=\int\mathrm{d}^{4}x\sqrt{-h_{4}}\left(-\sigma-\frac{\sigma\tau^{2}}{2}|\mathcal{F}|^{2}\right)$ (108)
This is the Maxwell action where $\sigma\tau^{2}$ is the inverse gauge coupling squared. This field strength sources the $H$-field. Indeed, we find that